OpenStudy (anonymous):

Let f(x) = Ax + B and g(x) = Cx + D where A, B, C, and D are non-zero constants. Find the vertical intercept of f(g(x)). Do not write your answer as a point.

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

I have solved f(g(x)) = A(Cx+D)+B, and I assumed the vertical intercept would just be B. However, that was a wrong answer.

OpenStudy (loser66):

Of course, it is wrong. hihihi

OpenStudy (loser66):

now open it : f(g(x))= A(Cx +D)+B= ACx +AD +B See it now??

OpenStudy (anonymous):

In that case, would you have to plug in 0 for x, which makes it: f(g(x)) = AD+B Making the vertical intercept: f(g(x)) - AD = B?

OpenStudy (loser66):

your plugging in is ok, but the conclusion is not good. It should be: "the vertical intercept of f(g(x)) is at y = AD+B

OpenStudy (anonymous):

Oh, so the way you conclude the problem matters. Thank you so much, that was the accepted answer!